Need help on this as soon as possible

Monark Corp. just paid an annual dividend of $1.50 and expects to pay a dividend of $1.80 next year, $2.20 in two years, and $2.55 in three years. Subsequently, Monark expects dividends to grow at a rate consistent with its expected earnings retention/investment rate of 60% and its expected realized return on equity of 15%. The market requires a return of 12% on Monark stock. What is your best estimate of Monark's current stock price and stock price three years from now?
Answer #1

According to Gordon Growth Model,

Terminal Value of Firm = Dividend expected to be paid next year / (Required Rate of Return - Growth Rate)

where, Growth rate (g) is the product of expected earnings retention (b) and expected realized return on equity (r)

Therefore, g = br

Hence, g = (0.60 * 15%) = 9%

Required Rate of Return = 12%

Attaching the Excel Sheet showing expected dividends, their discounted values and terminal value :-

A B Required Rate of Return Growth Rate 12% 9.00% Year Dividend PV @ 12% Discounted Values of Dividend 1.80 0.8928571 1.61 2.

AB Required Rate of Return 0.12 Growth Rate =15%*0.6 Year 1 1.8 2 2.2 3 2.55 Dividend PV @ 12% =1/(1+$D$3)^B7 =1/(1+$D$3)^B8

Adding up the Discounted values of Dividend with Terminal Value we get = $71.12

Hence, expected value of stock = $71.12

Year 1.61 1.75 1.82 65.95 71.12 SUM